An experiment involves studying the composition of genders of a set of triplets.what is probability that there are 2 boys and a girl in the sets

1
Expert's answer
2022-04-04T14:22:53-0400

"p=P\left( boy \right) =0.5\\X\,\,-\,\,the\,\,number\,\,of\,\,boys\,\,of\,\,3 kids, X\sim Bin\left( 3,0.5 \right) \\P\left( X=2 \right) =C_{3}^{2}\cdot 0.5^2\cdot \left( 1-0.5 \right) ^1=0.375"